Find the rate of change of the function f(x)= x^3 with respect to x at x0.

1
Expert's answer
2022-06-02T16:46:12-0400
"f'(x)=(x^3)'=3x^2"

The rate of change of the function "f(x)= x^3" with respect to "x" at "x_0" is "3x_0^2."